# Primordial Black Holes from Primordial Voids

### v1

@@ A.1 Evolution equations
$S_{a}$ $\equiv\gamma^{rr}S_{rr}=\frac{1}{ae^{4\chi}}(E+D+p)+p~,$ (45)
$S_{b}$ $\equiv\gamma^{\theta}S_{\theta\theta}=r^{2}\frac{b}{a}~.$ (46)
@@
$\>-D_{k}\left[\alpha v^{k}\left(\mathcal{E}+D\right)\right]+\alpha K(\mathcal{E}+p)~,$ (49)
where $D_{i}$ denotes the covariant derivative with respect to the spatial metric. A generic equation of state often requires a root-finding technique to obtain the value of $p$ and $v^{r}$, prior to the recovery of the other variables, as they are needed to evolve Eqs. (47)–(49). However, for cases with a simple equation of state, the recovery of $p$ and $v^{r}$ can be done analytically by finding the physical root of a high-order polynomial. For a barotropic fluid with $p=\omega\rho_{\rm fl}$, this implies solving a second-order polynomial Staelens et al. (2021), which diminishes the computational burden of the simulations.
@@ A.3 Construction of initial data
$S_{i}$ $=0~.$ (55)

### v2

@@ A.1 Evolution equations
$S_{a}$ $\equiv\gamma^{rr}S_{rr}=\frac{1}{ae^{4\chi}}(E+D+p)v_{r}^{2}+p~,$ (45)
$S_{b}$ $\equiv\gamma^{\theta\theta}S_{\theta\theta}=p~.$ (46)
@@
$\>-D_{k}\left[\alpha v^{k}\left(\mathcal{E}+p\right)\right]+\alpha K(\mathcal{E}+p)~,$ (49)
where $D_{i}$ denotes the covariant derivative with respect to the spatial metric and $\mathcal{L}_{\beta}$ the Lie derivative with respect to $\beta^{i}$. A generic equation of state often requires a root-finding technique to obtain the value of $p$ and $v^{r}$, prior to the recovery of the other variables, as they are needed to evolve Eqs. (47)–(49). However, for cases with a simple equation of state, the recovery of $p$ and $v^{r}$ can be done analytically by finding the physical root of a high-order polynomial. For a barotropic fluid with $p=\omega\rho_{\rm fl}$, this implies solving a second-order polynomial Staelens et al. (2021), which diminishes the computational burden of the simulations.
Figure 5: Formation and identification of the PBH’s apparent horizon for $\mu=0.8$ (top panels) and $\mu=-1.8$ (bottom panels) by using the $2M_{\rm MS}/R=1$ (left panels) and the $\Theta_{+}<0$ conditions (right panels). The vertical dashed black line marks the location of the AH, shortly after its formation (depicted by red solid curves).
Figure 6: Violation of the relative Hamiltonian constraint. Left panel shows local values at specific time throughout the simulations. Right panel shows the grid-averaged values, where the vertical color lines denotes the timing of the configuration showed in the left panel.
@@ A.3 Construction of initial data
$S_{i}$ $=0~,$ (55)
where the background quantities are set to follow the FLRW relations of a perfect fluid
$H_{\rm bkg}(t)=\frac{2}{3(1+\omega)t}~,\qquad\rho_{\rm bkg}(t)=\frac{3H_{\rm bkg}^{2}(t)}{8\pi}~.$ (56)
@@
A.4 Mass and Apparent Horizons
As described in section II.2, we can use both the Misner-Sharp relation (12) and the outgoing null expansion $\Theta_{+}$ to determine the formation of the PBH apparent horizon. Using the BSSN system describe above, these quantities read
$\Theta_{+}$ $=\frac{e^{-2\chi}}{\sqrt{a}}\left(4\,\partial_{r}\chi+\frac{2}{r}+\frac{\partial_{r}b}{b}\right)+A_{a}-\frac{2}{3}K~,$ (57)
$M_{\rm MS}$ $=\frac{1}{2}R\left[1+\left(\frac{\partial_{t}R-\beta^{r}\partial_{r}R}{\alpha}\right)^{2}-\frac{e^{-4\chi}}{a}(\partial_{r}R)^{2}\right]~,$ (58)
where
$\partial_{t}R$ $=e^{2\chi}r\sqrt{b}\left(2\,\partial_{t}\chi+\frac{1}{2}\frac{\partial_{t}b}{b}\right)~,$ (59)
$\partial_{r}R$ $=e^{2\chi}\sqrt{b}\left(1+2r\,\partial_{r}\chi+\frac{r}{2}\frac{\partial_{r}b}{b}\right)~.$ (60)
Figure 6 shows two examples of the AH formation and identification using these quantities for the cases $\mu=-1.80$, $\mu=-0.8$ with $\omega=1/3$.
A.5 Code validation
As customary, we check for the validity of our simulations by checking the fulfillment of the constraint equations (50-51). In particular, we use that relative violation of the Hamiltonian constraint be small, at least $\left|\mathcal{H}_{\rm rel}<10^{-2}\right|$, where
$\mathcal{H}_{\rm rel}\equiv~\frac{{}^{(3)}R+K^{2}-K_{ij}K^{ij}-16\pi\rho}{|^{(3)}R|+|K^{2}|+|K_{ij}K^{ij}|+|16\pi\rho|}~.$ (61)
Larger violations deep inside the AH are ignored (near $r\sim 0$), as long stability outside the AH holds.
Figure 6   shows an example of code validation for the case $\mu=-1.80$ with $\omega=1/3$.